Circular No. 7395

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) G. H. Jones and A. Balogh, Imperial College, London; and T. S. Hor-

bury, Queen Mary and Westfield College, London, write: “Analysis of magnetometer data returned from the ESA/NASA Ulysses spacecraft for several hours centered on 1996 May 1.375 UT reveals solar-wind magnetic-field signatures consistent with a cometary ion-tail crossing. Our search for a source comet revealed that C/1996 B2 passed across the sun-Ulysses line during 1996 Apr. 23. This 8-day time difference equals the travel time for cometary ions to be carried anti-sunward to Ulysses at up to 750 km/s (the exact travel time depending on the acceleration rate within the tail). Magnetic-field vectors were consistent with a tail parallel to the comet’s orbital plane, as expected, and the tail diameter of > 7 × 10^{6} km (inferred from the magnetometer data) is consistent with that expected for gas-production-rate estimates for the comet on Apr. 23. Consideration of the relative positions of Ulysses and the comet’s nucleus implies a tail length > 3.8 AU. Details are given by Jones et al. (2000, Nature 404, 574). We also note that the earth-comet-sun phase angle during the Ulysses tail crossing on May 1 was 32^{o}, and if the tail (out to the position of Ulysses) had been then visible from the earth (with the comet’s head at solar elongation ∼ 7^{o}), its apparent length would have been > 80^{o}. This is due to the tail’s curvature, which is clear in the magnetometer data. This curvature may provide an explanation for the reported apparent tail lengths in late Mar. 1996 that were larger than the phase angle (cf. IAUC 6360).”

XTE J1118+480 V. Dhawan, National Radio Astronomy Observatory; G. G. Pooley,

Mullard Radio Astronomy Observatory, University of Cambridge; and R. N. Ogley and I. F. Mirabel, Service d’Astrophysique, Laboratoire du Commissariat à l’Energie Atomique, Saclay, report on additional radio photometry of XTE J1118+480 (cf. IAUC 7390) using the Ryle and VLA telescopes: “Observations at 15 GHz with the Ryle telescope on Mar. 31.77–31.89 and Apr. 2.74–02.88 UT measure flux densities of 7.8 ± 0.35 and 7.5 ± 0.30 mJy. Observations with the VLA on Apr. 2.98–3.13 produce flux densities at 1.4, 8.3, and 23 GHz of 2.1, 6.0 ± 0.1, and 8.8 ± 0.3 mJy. The position of the source measured by the VLA at 23 GHz is α = 11^{h}18^{m}10^{s}.79, δ = +48^{o}02′12′′.3 (equinox 2000.0; uncertainty 0′′.3). The inverted radio spectrum is consistent with emission from an optically thick ejection.”

2000 April 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green